Find the inverse of the matrix
[6 7]..........................
[2 3]
Use this inverse to find the coordinates of the point of intersection of the line 5c+7y=19 and 2x+3y=8 Explain why this method fail when applied to the lines 3x+4y=7 and 6x+8y=9

Answer by ikleyn(52775)

.
Find the inverse of the matrix
[6 7]..........................
[2 3]
Use this inverse to find the coordinates of the point of intersection of the line 5c+7y=19 and 2x+3y=8
Explain why this method fail when applied to the lines 3x+4y=7 and 6x+8y=9
~~~~~~~~~~~~~~~


Your matrix to inverse  DOES  NOT  CORRESPOND  to the system of equation in this post,
which you want to solve.

Extremely inaccurate writing/posting.